# black holes; early universe; general relativity

### v1

@@ 5.3 Type B horizon structure
To describe the horizons of the cosmological black holes, we introduce the trapped spheres and trapping horizons. See Refs. Hayward (1994, 1996) for the notion of trapped spheres and trapping horizons. For a spherically symmetric spacetime, we can generally introduce radial null coordinates $\xi_{\pm}$ such that the line element can be written in the following form:
$ds^{2}=-2f(\xi_{+},\xi_{-})d\xi_{+}d\xi_{-}+R^{2}(\xi_{+},\xi_{-})(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (24)
We introduce the future-directed radial null vectors $l^{a}_{\pm}\propto(\partial_{\pm})^{a}$ such that $g_{ab}l^{a}_{+}l^{b}_{-}=-1$, where ±$\partial_{\pm}=\partial/\partial\xi_{\pm}$. Then, we call $\theta_{+}$ and $\theta_{-}$ outgoing and ingoing null expansions, respectively, where
$\theta_{\pm}:=l^{a}_{\pm}\partial_{a}\ln(R^{2}).$ (25)
We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a future (past) trapped sphere if $\theta_{+}\theta_{-}>0$ and $\theta_{+}+\theta_{-}<(>)0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a future (past) marginal sphere if $\theta_{+}\theta_{-}=0$ and $\theta_{+}+\theta_{-}<(>)0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a bifurcating marginal sphere if $\theta_{+}=\theta_{-}=0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ an untrapped sphere if $\theta_{+}\theta_{-}<0$. We call a spacetime region a future (past) trapped region, if any 2-sphere given by $(\xi_{+},\xi_{-})$ in the region is a future (past) trapped sphere. We call a spacetime region an untrapped region, if any 2-sphere given by $(\xi_{+},\xi_{-})$ in the region is an untrapped sphere. We call a hypersurface foliated by future (past) marginal spheres a future (past) trapping horizon. We call a hypersurface foliated by bifurcating marginal spheres a bifurcating trapping horizon Maeda et al. (2009).
@@
Figure 9: Structures of trapping horizons and trapped regions inferred by numerical simulations. The structures indicated on the top, middle and bottom panels are called type A, marginal and type B, respectively. The signs of the radial null expansions $({\rm sign}(\theta_{-}),{\rm sign}(\theta_{+}))$ are shown for each region. The regions are divided by trapping horizons, where $\theta_{+}\theta_{-}=0$. The future and past trapping horizons are denoted by the red and blue curves, respectively. The bifurcating trapping horizons are denoted by open circles. The black dashed lines denote spacetime singularities, while the black solid lines denote the regular centres and the null infinities. This figure is comparable to Fig. 7 of Ref. Uehara et al. (2024).
@@
With the above terminology, we can now describe the structure of trapping horizons. We specify $\xi_{+}$ and $\xi_{-}$ as the null coordinates going to the outgoing and ingoing null coordinates in the far region which is asymptotic to the flat FLRW universe, which we assume to exist. There is a big difference in the structure of trapping horizons and trapped regions between $\mu=1.2$ and $\mu=1.8$. We schematically plot the horizon structures inferred by numerical simulations in Fig. 9. For $\mu=1.2$ as shown on the left panel, there are a past trapped region and a future-trapped region that are disconnected from each other with the associated past and future trapping horizons being separate. We call this horizon structure type A. For $\mu=1.8$ as shown on the right panel, it is very different. The past and future trapped regions have contacts at two points. The two points of intersection, which are in fact two 2-spheres, correspond to bifurcating trapping horizons. There appears an untrapped region in the centre enclosed by the future and past trapping horizons and the two bifurcating trapping horizons, where $\theta_{+}<0$ and $\theta_{-}>0$. We call this horizon structure type B.
@@
This numerical result for the perturbation indicates that there appears the peculiar structure of trapping horizons and trapped regions, if $\mu$ is greater than some critical value. This was also true if we took another choice of $\zeta(r)$, so the above features are at least general to some extent. So, we can regard this structure as common for the sufficiently large amplitude of curvature perturbation of type II. We infer the existence of the marginal horizon structure between types A and B, which is shown on the middle panel of Fig, 9. So, we can say that a type II perturbation does not always result in a type B structure for radiation. This is due to the effect of pressure because it is known that for the dust case a type II perturbation necessarily entails the horizon structure of type B. So, we can conclude that there are at least PBHs with structure of types IA, IIA and IIB in radiation domination.
6 Summary
@@
In the recent development of research on this subject, it has been revealed that PBHs lie at the intersection of various developing branches of modern physics.
@@
In Sec. 2, we present the basic concept of PBHs. The mass of PBHs is usually considered to be on the mass scale within the cosmological horizon at the time of their formation. Mass accretion may significantly increase the PBH mass depending on the evolution scenario, whereas it has been shown to be negligible during the evolution in radiation domination. The mass of PBHs can be significantly reduced by Hawking evaporation, and they would have evaporated completely by now if they were lighter than the critical mass of approximately $10^{15}$ g, although the details of the evaporation process are still somewhat under debate. We obtain constraints on the fraction $f(M)$ of PBHs of mass $M$ in relation to all dark matter through different observations. The fraction $f(M)$ can be transformed into the formation probability $\beta(M)$ depending on the cosmological evolutionary scenario. The standard cosmic history implies that a very small value of $\beta(M)$, as small as $\sim 10^{-17}$ for $M\sim 10^{17}$ g, can yield $f(M)=O(1)$, i.e., can explain all dark matter, because the PBHs’ contribution to the energy of the Universe increases in proportion to the scale factor during the radiation-dominated phase.
@@
In Sec. 3, we discuss PBH formation. A detailed and precise understanding of PBH formation physics has become increasingly important. The basic question in this study is how to predict $\beta(M)$ and other observationally significant quantities from a given cosmological scenario. Focusing on PBH formation from fluctuations generated by inflation, the key terms are inflation models, long-wavelength solutions, thresholds, softer EOS, matter domination, critical behaviours, and statistics.
@@
In Sec 4, we discuss the initial spins of PBHs. PBHs formed during radiation domination are unlikely to have large spins. Perturbative studies show that the non-dimensional Kerr parameter of these PBHs is typically of the order of $10^{-3}$. In contrast, PBHs formed during matter domination can acquire large spins, at least initially. The effect of mass accretion after formation on the non-dimensional Kerr parameter needs to be studied carefully. It is evident that numerical simulations based on numerical relativity should shed light on this problem.
@@
In Sec. 5, we review the recently implemented numerical relativity simulations of the time evolution of type II perturbations during radiation domination. This is particularly relevant to a very rare perturbation peak or to a specific inflationary scenario. The resulting structures of trapping horizons can be classified into two types: one is standard for PBHs, and the other is very unique, featuring the crossing of trapping horizons as bifurcating trapping horizons. We refer to these as types A and B, respectively, while we can also discuss the marginal structure. The numerical simulations suggest that the evolution of type II perturbations can be classified into type A and type B based on their horizon structure. Thus, we may call them types IIA and IIB.
@@
Finally, the author must acknowledge that there are many interesting issues concerning the formation of PBHs that cannot even be mentioned in this article. The study of PBH formation not only requires a deep understanding of physical phenomena within known standard physics but also offers the opportunity to explore unknown new physics through PBHs. Both of these aspects will play important roles in the future of PBH formation studies.

### v3

@@ 5.4 Type B horizon structure
In the asymptotically flat spacetimes, black holes are conventionally defined by event horizons. However, in cosmological setting, event horizons are not necessarily useful because the definition of an event horizon is teleological in the sense that it is only based on the infinite future and because the structure of infinities in cosmological spacetimes can be very different from asymptotically flat ones. Furthermore, in the expanding Universe, there may be another type of horizons, i.e., a cosmological horizon. As we will see below, we encounter the interplay of these two types of horizons for some class of PBH formation from the type II configuration. This implies that we need to distinguish between these two types of horizons. For these reasons, we adopt the notion of trapping horizons, a quasilocal formulation of horizons Hayward (1994, 1996). This is akin to the notion of apparent horizons in the present setting. Here, we define horizons using the pair of null expansions $\theta_{+}$ and $\theta_{-}$ along outgoing and ingoing radial null coordinates $\xi_{+}$ and $\xi_{-}$, respectively. Black hole horizons can be defined with one vanishing and one negative null expansion so that any radial light ray there cannot expand its cross section. Similarly, cosmological horizons can be defined with one vanishing and one positive null expansion so that any radial light ray there cannot contract its cross section. The black hole horizons and cosmological horizons are called future trapping horizons and past trapping horizons, respectively. In principle, we may also have a horizon with both null expansions vanishing, which is called a bifurcating trapping horizon Maeda et al. (2009). In Appendix A, we will introduce trapping horizons in a more mathematical manner, whereas the intuitive understanding discussed above is still useful.
@@
Figure 10: Structures of trapping horizons and trapped regions inferred by numerical simulations. The structures indicated on the top, middle and bottom panels are called type A, marginal and type B, respectively. The signs of the radial null expansions $({\rm sign}(\theta_{-}),{\rm sign}(\theta_{+}))$ are shown for each region. The regions are divided by trapping horizons, where $\theta_{+}\theta_{-}=0$. The future and past trapping horizons are denoted by the red and blue curves, respectively. The bifurcating trapping horizons are denoted by open circles. The black dashed lines denote spacetime singularities, while the black solid lines denote the regular centres and the null infinities. This figure is comparable to Fig. 7 of Ref. Uehara et al. (2024). See text for more details.
@@
With the above terminology, we can now describe the structure of trapping horizons. We assign $\xi_{+}$ and $\xi_{-}$ as the null coordinates being the standard ones in the far region which is asymptotic to the flat FLRW solution, which we assume to exist. There is a big difference in the structure of trapping horizons and trapped regions between $\mu=1.2$ and $\mu=1.8$. We schematically plot the horizon structures inferred by numerical simulations in Fig. 10. For $\mu=1.2$, as shown on the top panel, there are a past trapped region and a future-trapped region that are disconnected from each other with the associated past and future trapping horizons being separate. We call this horizon structure type A. For $\mu=1.8$, as shown on the bottom panel, the past and future trapped regions have contacts at two points. The two points of intersection, which are in fact two 2-spheres, correspond to bifurcating trapping horizons. There appears an untrapped region enclosed by the future and past trapping horizons and the two bifurcating trapping horizons, where $\theta_{+}<0$ and $\theta_{-}>0$. We call this horizon structure type B.
@@
This numerical result for the perturbation indicates that there appears the peculiar structure of trapping horizons and trapped regions, if $\mu$ is greater than some critical value. This was also true if we took another choice of $\zeta(r)$, so the above features are at least general to some extent. So, we can regard this structure as common for a sufficiently large amplitude of curvature perturbation of type II. We infer the existence of the marginal horizon structure between types A and B, which is shown on the middle panel of Fig. 10. We can say that a type II perturbation does not always result in a type B structure for radiation. This is due to the effect of pressure because it is known that for the dust case a type II perturbation necessarily entails the horizon structure of type B. So, we can conclude that there are at least PBHs with structure of types I-A, II-A and II-B in radiation domination.
It should be noted that the type B horizon structure has not been known in the gravitational collapse in asymptotically flat spacetimes and therefore very unique to PBH formation as long as the author is aware. For this reason, the result of the type II-B PBH formation is very intriguing in the context of the variety of black hole formation in general relativity. On the other hand, their relevance to observational cosmology is yet unclear.
6 Conclusions
@@
In the recent development of research on this subject, it has been revealed that PBHs lie at the intersection of various developing branches of modern physics. In this article, we review their basic concept, formation, spins and link to the type II configuration.
@@
In Sec. 2, we present the basic concept of PBHs. The mass of PBHs is usually considered to be the mass scale within the cosmological horizon at the time of their formation. Mass accretion may significantly increase the PBH mass depending on the evolution scenario, whereas it has been shown to be negligible during the evolution in radiation domination. The mass of PBHs can be significantly reduced by Hawking evaporation, and they would have evaporated completely by now if they were lighter than the critical mass of approximately $10^{15}$ g, although the details of the evaporation process are still somewhat under debate. We can obtain constraints on the fraction $f(M)$ of PBHs of mass $M$ in relation to all dark matter through different observations. The fraction $f(M)$ can be transformed into the formation probability $\beta(M)$ depending on the cosmological evolutionary scenario. The standard cosmic history implies that a very small value of $\beta(M)$, as small as $\sim 10^{-17}$ for $M\sim 10^{17}$ g, can yield $f(M)=O(1)$, i.e., can explain all dark matter, because the PBHs’ contribution to the energy of the Universe increases in proportion to the scale factor during the radiation-dominated phase.
@@
In Sec. 3, we discuss PBH formation. A detailed and precise understanding of PBH formation physics has become increasingly important. The basic question in this study is how to predict $\beta(M)$ and other observationally significant quantities from a given cosmological scenario. Focusing on PBH formation from fluctuations generated by inflation, the key terms are inflation models, long-wavelength solutions, thresholds, softer EOS, matter domination, critical behaviours and statistics.
@@
In Sec 4, we discuss the initial spins of PBHs. PBHs formed during radiation domination are unlikely to have large spins. Perturbative studies show that the nondimensional Kerr parameter of these PBHs is typically of the order of $10^{-3}$. In contrast, PBHs formed during matter domination can acquire large spins, at least initially. The effect of mass accretion after formation on the nondimensional Kerr parameter needs to be studied carefully. It is evident that numerical simulations based on numerical relativity should shed light on this problem.
@@
In Sec. 5, we review the recently conducted numerical relativity simulations of the time evolution of type II perturbations during radiation domination. This is particularly relevant to a very rare perturbation peak or to a specific inflationary scenario. The resulting structures of trapping horizons can be classified into two types: one is standard for PBHs, and the other is very unique, featured with the crossing of trapping horizons as bifurcating trapping horizons. We refer to these as types A and B, respectively, while we can also discuss the marginal structure. The numerical simulations suggest that the evolution of type II perturbations can be classified into type A and type B based on their horizon structure. Thus, we may call them types II-A and II-B.
@@
Finally, I must acknowledge that there are many interesting issues concerning the formation of PBHs that cannot even be mentioned in this article. The study of PBH formation not only requires a deep understanding of physical phenomena within known standard physics but also offers the opportunity to explore unknown new physics through PBHs. Both of these aspects will play important roles in the future of PBH formation studies.
@@ Acknowledgements.
Appendix A Trapping horizons
For a spherically symmetric spacetime, we can generally introduce radial null coordinates $\xi_{\pm}$ such that the line element can be written in the following double-null form:
$ds^{2}=-2e^{-f(\xi_{+},\xi_{-})}(\xi_{+},\xi_{-})d\xi_{+}d\xi_{-}+R^{2}(\xi_{+},\xi_{-})(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (31)
We introduce the future-directed radial null vectors $l^{a}_{\pm}\propto(\partial_{\pm})^{a}$ such that $g_{ab}l^{a}_{+}l^{b}_{-}=-1$, where ±$\partial_{\pm}=\partial/\partial\xi_{\pm}$. Then, we define $\theta_{\pm}$ as
$\theta_{\pm}:=l^{a}_{\pm}\partial_{a}\ln(R^{2}).$ (32)
We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a future (past) trapped sphere if $\theta_{+}\theta_{-}>0$ and $\theta_{+}+\theta_{-}<(>)0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a future (past) marginal sphere if $\theta_{+}\theta_{-}=0$ and $\theta_{+}+\theta_{-}<(>)0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ a bifurcating marginal sphere if $\theta_{+}=\theta_{-}=0$. We call a 2-sphere specified with $(\xi_{+},\xi_{-})$ an untrapped sphere if $\theta_{+}\theta_{-}<0$. We call a spacetime region a future (past) trapped region, if any 2-sphere given by $(\xi_{+},\xi_{-})$ in the region is a future (past) trapped sphere. We call a spacetime region an untrapped region, if any 2-sphere given by $(\xi_{+},\xi_{-})$ in the region is an untrapped sphere. We call a hypersurface foliated by future (past) marginal spheres a future (past) trapping horizon. We call a hypersurface foliated by bifurcating marginal spheres a bifurcating trapping horizon.
See Refs. Hayward (1994, 1996); Maeda et al. (2009) for more complete rigorous discussions and proofs of the basic properties of trapping horizons.